# Holistic analysis on the sustainability of Federated Learning across AI product lifecycle

### v2

@@ Abstract
In light of emerging legal requirements and policies focused on privacy protection, there is a growing trend of companies across various industries adopting Federated Learning (FL). This decentralized approach involves multiple clients or silos, collaboratively training a global model under the coordination of a central server while utilizing their private local data. Unlike traditional methods that necessitate data sharing and transmission, Cross-Silo FL allows clients to share model updates rather than raw data, thereby enhancing privacy. Despite its growing adoption, the carbon impact associated with Cross-Silo FL remains poorly understood due to the limited research in this area. This study seeks to bridge this gap by evaluating the sustainability of Cross-Silo FL throughout the entire AI product lifecycle, extending the analysis beyond the model training phase alone. We systematically compare this decentralized method with traditional centralized approaches and present a robust quantitative framework for assessing the costs and CO2 emissions in real-world Cross-Silo FL environments. Our findings indicate that the energy consumption and costs of model training are comparable between Cross-Silo Federated Learning and Centralized Learning. However, the additional data transfer and storage requirements inherent in Centralized Learning can result in significant, often overlooked CO2 emissions. Moreover, we introduce an innovative data and application management system that integrates Cross-Silo FL and analytics, aiming at improving the sustainability and economic efficiency of IT enterprises.
Index Terms:
Sustainable AI, Carbon Emission Estimation, Cross-Silo Federated Learning, Neural Networks
I Introduction
@@
The Information and Communication Technology (ICT) sector has experienced substantial and rapid growth over the past few decades, contributing significantly to global CO2 emissions [1]. This upward trajectory is expected to accelerate, fueled by the increasing reliance on technology such as Artificial Intelligence (AI) in everyday activities [1, 2]. Innovations in Machine Learning (ML) and Deep Learning (DL) have greatly enhanced efficiency and accuracy across various sectors. However, these advancements have also resulted in the development of more complex models with larger parameter sizes and larger training data size, necessitating considerable energy consumption and substantial water usage for cooling data centers [3]. For example, the training of the GPT-3 model which comprises 175 billion parameters reportedly consumed around 1,287 MWh of electricity, leading to carbon emissions estimated at approximately 500 metric tons of CO2 [4, 5, 6]. To contextualize this, the energy used is sufficient to power about 121 average households in the United States for a full year [3]. Beyond the training phase, other components of the AI product lifecycle, including data storage, model inference, and model management, also have substantial environmental impacts. Consequently, it is imperative to conduct comprehensive studies on the carbon footprint associated with the entire lifecycle of AI products. Furthermore, the development and implementation of effective strategies to enhance sustainability within the ICT sector are essential.
@@
In traditional AI model training, data is usually aggregated in a centralized manner, a process known as Centralized Learning (CL). However, CL presents significant vulnerabilities as centralized data repositories are prone to cyber-attacks, potentially leading to grave repercussions. Furthermore, when the data encompass personal or proprietary information, it also raises profound privacy concerns and complicates the collection and training processes within centralized systems [7]. To mitigate these privacy issues, Federated Learning (FL) [8] has emerged as an innovative solution. FL is a distributed learning framework that emphasizes privacy preservation by allowing multiple clients to collaboratively train a global model under the guidance of a central server [9]. Instead of sharing raw data, clients utilize their private local datasets to train models and subsequently share the model updates. When the clients participating in the training are small or distributed entities with limited local data, the process is known as cross-device Federated Learning. On the other hand, Cross-Silo Federated Learning involves a smaller number of clients, typically organizations or companies, each possessing substantial local datasets [10].
@@
The paper is organized as follows: the background and related works are introduced in Section II. In Section III, the novel framework for the analysis of energy and carbon footprints in Cross-Silo FL is introduced. We describe the experiments comparing the CO2 emission and cost of Cross-Silo Federated Learning with Centralized Learning in Section IV and the proposed sustainable data and application management system in V. Finally, the conclusion and future works are given in Section VI.
II Related Works
@@
III Quantifying CO2 estimations
III-A The real world Federated Learning environment
@@ III-B CO2 estimation on the Cloud
Cloud providers typically offer storage, database services, and node pools that inherently replicate data along with associated computational and memory resources to ensure high data durability and availability. This built-in redundancy is crucial for maintaining service continuity during outages. For example, ”Azure Storage always stores multiple copies of your data so that it’s protected from planned and unplanned events, including transient hardware failures, network or power outages, and massive natural disasters” [^1] https://learn.microsoft.com/en-us/azure/storage/common/storage-redundancy. Locally redundant storage (LRS) achieves this by synchronously replicating your data three times within a single physical location in the primary region. On the other hand, zone-redundant storage (ZRS) synchronously replicates data across three different Azure availability zones within the primary region. It is essential to account for data redundancy when evaluating the carbon footprint associated with data transfer and storage.
@@ IV Experiments
IV-A Datasets and experimental protocols
@@ IV-B Experimental results
V Proposed Data and application management system
Figure 7: The proposed Sustainable Data and Application management system leveraging Cross-Silo Federated Learning and Federated Analytics.
@@ V Proposed Data and application management system
  • The application or product owner will then employ Cross-Silo Federated Learning or Federated Analytics for further development. This method ensures efficient use of data and computational resources, fostering a more sustainable and collaborative environment for data and application management within the enterprise.
VI Conclusion

### v3

# Holistic analysis on the sustainability of Federated Learning lifecycle in real-world industrial settings

@@ Abstract
In light of emerging legal requirements and policies focused on privacy protection, there is a growing trend of companies across various industries adopting Federated Learning (FL). This decentralized approach involves multiple clients or silos, collaboratively training a global model under the coordination of a central server while utilizing their private local data. Unlike traditional methods that require data sharing and transmission, Cross-Silo FL allows clients to share model updates rather than raw data, thereby enhancing privacy. Despite its growing adoption, the carbon impact associated with Cross-Silo FL remains poorly understood due to the limited research in this area. Furthermore, most existing studies rely on simulated environments rather than real-world scenarios. This study seeks to bridge this gap by evaluating the sustainability of Cross-Silo FL throughout the entire AI product lifecycle in real-world industrial settings, extending the analysis beyond the model training phase alone. We systematically compare this decentralized method with traditional centralized approaches and present a robust quantitative framework for assessing the costs and CO2 emissions in real-world Cross-Silo FL environments. Our findings indicate that the energy consumption and costs of model training are comparable between Cross-Silo Federated Learning and Centralized Learning. However, the additional data transfer and storage requirements inherent in Centralized Learning can result in significant, often overlooked CO2 emissions. Moreover, we introduce an innovative data and application management system that integrates Cross-Silo FL and analytics, aimed at improving the sustainability and economic efficiency of IT enterprises. This study highlights the real-world applicability and benefits of Federated Learning, providing valuable insights and lessons learned that can be leveraged to enhance the deployment of AI methods in industrial applications.
[…]
123
1 Introduction
@@
The Information and Communication Technology (ICT) sector has experienced substantial and rapid growth over the past few decades, contributing significantly to global CO2 emissions [9]. This upward trajectory is expected to accelerate, fueled by the increasing reliance on technology such as Artificial Intelligence (AI) in everyday activities [9, 25, 13]. Innovations in Machine Learning (ML) and Deep Learning (DL) have greatly enhanced efficiency and accuracy across various sectors [17, 16, 14]. However, these advancements have also resulted in the development of more complex models with larger parameter sizes and larger training data size, necessitating considerable energy consumption and substantial water usage for cooling data centers [10]. For example, the training of the GPT-3 model which comprises 175 billion parameters reportedly consumed around 1,287 MWh of electricity, leading to carbon emissions estimated at approximately 500 metric tons of CO2 [29, 21, 24]. To contextualize this, the energy used is sufficient to power about 121 average households in the United States for a full year [10]. Beyond the training phase, other components of the AI product lifecycle, including data storage, model inference, and model management, also have substantial environmental impacts. Consequently, it is imperative to conduct comprehensive studies on the carbon footprint associated with the entire lifecycle of AI products. Furthermore, the development and implementation of effective strategies to enhance sustainability within the ICT sector are essential.
@@
In traditional AI model training, data is usually aggregated in a centralized manner, a process known as Centralized Learning (CL). However, CL presents significant vulnerabilities as centralized data repositories are prone to cyber-attacks, potentially leading to grave repercussions. Furthermore, when the data encompass personal or proprietary information, it also raises profound privacy concerns and complicates the collection and training processes within centralized systems [32]. To mitigate these privacy issues, Federated Learning (FL) [28] has emerged as an innovative solution. FL is a distributed learning framework that emphasizes privacy preservation by allowing multiple clients to collaboratively train a global model under the guidance of a central server [20, 19]. Instead of sharing raw data, clients utilize their private local datasets to train models and subsequently share the model updates. When the clients participating in the training are small or distributed entities with limited local data, the process is known as cross-device Federated Learning. On the other hand, Cross-Silo Federated Learning involves a smaller number of clients, typically organizations or companies, each possessing substantial local datasets [30].
@@
The paper is organized as follows: the background and related works are introduced in Section 2. In Section 3, the novel framework for the analysis of energy and carbon footprints in Cross-Silo FL is introduced. We describe the experiments comparing the CO2 emission and cost of Cross-Silo Federated Learning with Centralized Learning in Section 4 and the proposed sustainable data and application management system in Section 5. Finally, the conclusion and future works are given in Section 6.
2 Related Works
@@
3 Quantifying CO2 estimations
3.1 The real world Federated Learning environment
@@ 3.2 CO2 estimation on the Cloud
Cloud providers typically offer storage, database services, and node pools that inherently replicate data along with associated computational and memory resources to ensure high data durability and availability. This built-in redundancy is crucial for maintaining service continuity during outages. For example, "Azure Storage always stores multiple copies of your data so that it’s protected from planned and unplanned events, including transient hardware failures, network or power outages, and massive natural disasters" [^1] https://learn.microsoft.com/en-us/azure/storage/common/storage-redundancy. Locally redundant storage (LRS) achieves this by synchronously replicating your data three times within a single physical location in the primary region. On the other hand, zone-redundant storage (ZRS) synchronously replicates data across three different Azure availability zones within the primary region. It is essential to account for data redundancy when evaluating the carbon footprint associated with data transfer and storage.
@@ 4 Experiments
4.1 Datasets and experimental protocols
@@ 4.2 Experimental results
5 Proposed Data and application management system
Figure 7: The proposed Sustainable Data and Application management system leveraging Cross-Silo Federated Learning and Federated Analytics: Application owners submit use cases to data owners, which are checked for redundancy using historical requests and LLMs. Approved requests are granted data access and compute resources. FL and FA then enable privacy-preserving, resource-efficient development across organizational data silos.
@@ 5 Proposed Data and application management system
  • The application or product owner will then employ Cross-Silo Federated Learning or Federated Analytics for further development (solutions in the multi-view learning literature such as [15, 18, 12, 27] can be used too). This method ensures efficient use of data and computational resources, fostering a more sustainable and collaborative environment for data and application management within the enterprise.
6 Conclusion